# CEval: A Benchmark for Evaluating Counterfactual Text Generation

### v1

@@ Abstract
Counterfactual text generation aims to minimally change a text, such that it is classified differently. Judging advancements in method development for counterfactual text generation is hindered by a non-uniform usage of data sets and metrics in related work. We propose CEval, a benchmark for comparing counterfactual text generation methods. CEval unifies counterfactual and text quality metrics, includes common counterfactual datasets with human annotations, standard baselines (MICE, GDBA, CREST) and the open-source language model LLAMA-2. Our experiments found no perfect method for generating counterfactual text. Methods that excel at counterfactual metrics often produce lower-quality text while LLMs with simple prompts generate high-quality text but struggle with counterfactual criteria. By making CEval available as an open-source Python library, we encourage the community to contribute more methods and maintain consistent evaluation in future work. [^1] The source code is included with the paper submission and will be publicly accessible upon acceptance.
@@
Van Bach Nguyen University of Marburg / Germany vanbach.nguyen@uni-marburg.de                        Jörg Schlötterer University of Marburg / Germany University of Mannheim / Germany joerg.schloetterer@uni-marburg.de
Christin Seifert University of Marburg / Germany christin.seifer@uni-marburg.de
@@ 1 Introduction
The growing popularity of AI and increasingly complex “black-box” models triggered a critical need for interpretability. As Miller (2019) highlights, explanations are often counterfactual, seeking to understand why event P occurred instead of alternative Q. Ideally, explanations should demonstrate how minimal changes in an instance could have led to different outcomes. For example, given a review: The film has funny moments and talented actors, but it feels long. To answer the question why this review has a negative sentiment instead of a positive sentiment, the answer might involve showing a positive example as a counterfactual like: The film has funny moments and talented actors, yet feels a bit long. This example enables us to identify specific words that require change and the necessary modifications to achieve the target sentiment. This motivates counterfactual generation, the task of modifying an instance to produce a different model prediction with minimal change.
Besides explanation Robeer et al. (2021), the NLP community also utilizes counterfactual generation for various purposes such as debugging models Ross et al. (2021), data augmentation Dixit et al. (2022); Chen et al. (2023); Bhattacharjee et al. (2024) or enhancing model robustness Treviso et al. (2023); Wu et al. (2021). However, generating counterfactuals is not straightforward due to the complexity of textual changes, involving replacements, deletions, and insertions. Determining where and how to modify text to alter predictions remains an open issue. Existing research efforts unified evaluation standards and often prioritize quantity over quality in generated counterfactuals. Table 1 illustrates the disparity in datasets, metrics, and baselines across different studies. Consequently, it becomes challenging to choose an optimal method for a specific purpose. This highlights the need for a well-defined benchmark to comprehensively evaluate counterfactual generation methods for textual data. Such a benchmark should establish standard datasets, metrics, and baselines, enabling fair and meaningful comparisons.
@@
GBDA Guo et al. (2021) AG News, Yelp IMDB, MNLI Accuracy, Cosine Similarity, #Queries TextFooler Jin et al. (2020) BAE Garg and Ramakrishnan (2020) BERT-Attack Li et al. (2020)
CF-GAN Robeer et al. (2021) HATESPEECH, SST-2, SNLI Fidelity Perceptibility Naturalness SEDC Martens and Provost (2014) PWWS+ Ren et al. (2019) Polyjuice Wu et al. (2021) TextFooler Jin et al. (2020)
Polyjuice Wu et al. (2021) IMDB, NLI, QQP Diversity, Closeness GPT-2 Radford et al. (2019) T5 Raffel et al. (2020) RoBERTa Liu et al. (2019)
Table 1: The variations in datasets, metrics, and baselines used across different methods.
This work introduces CEval - a comprehensive benchmark to unify the evaluation of methods that modify text to change a classifier’s prediction. Such methods include contrastive explanations, counterfactual generation, and adversarial attacks. Our benchmark offers multifaceted evaluation: assessing both the “counterfactual-ness” (e.g., label flipping ability) and textual quality (e.g., fluency, grammar, coherence) of counterfactual texts. Additionally, it comprises carefully curated datasets with human annotations, as well as a straightforward baseline generated using a large language model. Moreover, we systematically review state-of-the-art methods in this field and directly compare their performance in our benchmark. We provide an open-source library for both the benchmark and the methods, promoting reproducibility and facilitating further research in this domain. Furthermore, we present a systematic comparison of different methods, highlighting their strengths and weaknesses in generating counterfactual text. We analyze how automatically generated counterfactuals compare to human examples, revealing gaps and opportunities for improvement. We find that counterfactual generation methods often generate text that lacks in quality compared to simple prompt-based LLMs. While the latter may struggle to satisfy counterfactual metrics, they typically exhibit higher text quality. Based on these insights, we suggest exploring combinations of methods as promising directions for further research.
@@ 2 Related Work
While terms like “counterfactual” and “contrastive” generation are often used interchangeably in literature Stepin et al. (2021), our work adopts a specific definition. We define counterfactual generation as generating a new instance with different model predictions from the original with minimal changes, encompassing both true counterfactuals and contrastives. This broader category includes counterfactual, contrastive generation, and adversarial attacks. In the past, adversarial attacks focused on changing the label without considering text quality. Recent work like GBDA Guo et al. (2021) focuses on producing adverserial text that is more natural by adding fluency and semantic similarity losses. Hence, we include GBDA in our benchmark. Technically, counterfactual generation methods for text fall into three categories:
@@
Masking and Filling Methods (MF): These methods perform 2 steps: (1) identify important words for masking and use various techniques such as selecting words with the highest gradient or training a separate rationalizer for the masking process. (2) The masked words are then replaced using a pretrained language model with fill-in-blank ability. In step (1), MICE Ross et al. (2021) and AutoCAD Wen et al. (2022) use the gradient of the classifier. DoCoGen Calderon et al. (2022) masks all relevant terms in a domain, while CREST Treviso et al. (2023) trains a separate rationalizer, i.e., SPECTRA Guerreiro and Martins (2021). Then, all of them fine-tune T5 to fill in blanks but train in different ways. Polyjuice Wu et al. (2021) accepts text with masking and fine-tunes a Roberta-based model to fill in blanks using control codes.
@@
Conditional Distribution Methods (CD): Methods like GBDA Guo et al. (2021) and CF-GAN Robeer et al. (2021) train a conditional distribution for counterfactuals. The counterfactuals are obtained by sampling from this distribution based on a target label.
@@
Counterfactual Generation with Large Language Models: Recently, there has been a trend towards using Language Models (LLMs) for counterfactual generation. Approaches like CORE Dixit et al. (2022), DISCO Chen et al. (2023) and FLARE Bhattacharjee et al. (2024) optimize prompts fed into LLMs to generate the desired counterfactuals. This trend is driven by the versatile capabilities of LLMs in various tasks Maynez et al. (2023).
@@
Despite the diverse approaches employed in generating counterfactuals across various studies, the common objective remains to generate high-quality counterfactuals. However, these studies employ different metrics, baselines, and datasets, as illustrated in Table 1. Therefore, given the rapid growth of approaches in this field, establishing a unified evaluation standard becomes paramount. Existing benchmarks for counterfactual generation Pawelczyk et al. (2021); Moreira et al. (2022) focus exclusively on tabular data with properties that are orthogonal to text (e.g., continuous value ranges). Hence, we introduce CEval to fill this gap and provide a standard evaluation framework specifically tailored to textual counterfactual generation. Our benchmark unifies metrics of both, counterfactual criteria and text quality assessment, including datasets with human annotations and a simple baseline from a large language model.
@@ 3 Benchmark Design
We focus on counterfactual generation for textual data, which involves editing existing text with minimal modifications to produce new text that increases the probability of a predefined target label with respect to a black-box model. This process aims to generate a counterfactual, denoted as $x^{\prime}$, that alters the model’s predictions compared to the original text $x$.
Formally, given a fixed classifier $f$ and a dataset with $N$ samples $(x_{1},x_{2},\ldots,x_{N})$, $x=(z_{1},z_{2},\ldots,z_{n})$ represents a sequence of $n$ tokens. The original prediction is denoted as $f(x)=y$, while the counterfactual prediction is $y^{\prime}\neq y$. The counterfactual generation process is represented by a method $e:(z_{1},\ldots,z_{n})\mapsto(z^{\prime}_{1},\ldots,z^{\prime}_{m})$, ensuring that $f(e(x))=y^{\prime}$. The resulting counterfactual example is $x^{\prime}=(z^{\prime}_{1},\ldots,z^{\prime}_{m})$ with $m$ tokens.
@@
Predictive Probability: A counterfactual instance $x^{\prime}$ should closely produce the predefined prediction $y^{\prime}$. In other words, the counterfactual text should effectively lead to the desired target label.
@@
Likelihood in Feature Space: A valid counterfactual instance should exhibit feature values that resemble real-world text, indicating that $x^{\prime}$ remains close to a common distribution for text. This distribution can be represented by pretrained large language models trained on massive text corpora. This criterion ensures that the generated text is plausible, realistic and consistent with typical language patterns.
Diversity: There are multiple ways to modify the input to reach the desired label. A good counterfactual method should provide various options for changing a text instance to obtain the target label.
@@ 3.1.1 Counterfactual metrics
Probability Change ($\Delta\textbf{P}$): While the flip rate offers a binary assesment of Predictive Probability, it does not fully capture how closely a counterfactual instance aligns with the desired prediction. Some instances may get really close to the target prediction but still fail to flip the label. To address this limitation, we propose the Probability Change ($\Delta$P) metric, which quantifies the difference between the probability of the target label for the original instance $x$ and the probability of the target label for the contrasting instance $x^{\prime}$.
$\Delta P=\frac{1}{N}\sum_{i=1}^{N}(P(y^{\prime}_{i}|x_{i},f)-P(y^{\prime}_{i}|x^{\prime}_{i},f))$
Here, $P(\cdot)$ represents the probability of having the prediction $y$ given the instance $x$ with respect to the classifier $f$. The expression denotes the difference in the probabilities of the prediction $y^{\prime}$ for instances $x$ and $x^{\prime}$ according to the classifier $f$. Token Distance (TD): To measure Textual Similarity, a common metric employed in literature Ross et al. (2021); Treviso et al. (2023) is the token-level distance $d(x,x^{\prime})$, where $d$ represents any edit distance metric. In this work, we use the Levenshtein distance for $d$ to quantify the token-distance between the original instance $x$ and the counterfactual $x^{\prime}$. This choice is motivated by the Levenshtein distance’s ability to capture all type of edits (insertions, deletions, or substitutions) and also its widespread use in related work Ross et al. (2021); Treviso et al. (2023):
$TD=\frac{1}{N}\sum_{i=1}^{N}\frac{{d(x_{i},x^{\prime}_{i}})}{|x_{i}|}$
Perplexity (PPL): To ensure that the generated text is plausible, realistic, and follows a natural text distribution, we leverage perplexity from GPT-2 because of its effectiveness in capturing such distributions. Radford et al. (2019) [^2] While GPT-2 is used for simplicity in this study, any other LLM can be substituted as long as it demonstrates strong text generation capabilities
@@
Diversity (Div): We quantify diversity by measuring the token distance between pairwise generated instances. Given two counterfactuals, $x^{\prime 1}$ and $x^{\prime 2}$, for the same instance $x$, diversity is defined as the mean pairwise distance between the sets of counterfactuals:
$Div=\frac{1}{N}\sum_{i=1}^{N}\frac{d(x^{\prime 1}_{i},x^{\prime 2}_{i})}{|x_{i}|}$
Here, $d(x^{\prime 1}_{i},x^{\prime 2}_{i})$ represents the Levenshtein distance between the corresponding tokens of the two counterfactuals for the $i$-th instance, and $|x_{i}|$ is the length of the original instance $x_{i}$.
@@ 3.1.2 Text Quality Metrics
In addition to counterfactual evaluation metrics, it is essential to consider the quality of the generated text. Text quality metrics are designed to evaluate specific aspects of textual content. Following Chiang and Lee (2023), key text quality metrics include:
Fluency: Relates to the natural and fluent flow of the generated text. A fluently constructed counterfactual instance should exhibit coherence in its structure and readability, ensuring that the transformed text reads naturally and seamlessly.
Cohesiveness: Focuses on the coherence and logical flow of ideas within the generated text. A cohesive counterfactual instance should deliver a cohesive narrative, with ideas and sentences seamlessly transitioning from one to another, contributing to an overall consistent and well-structured piece.
Likability: This metric focuses on how pleasant and likable the generated text is perceived to be by readers. It considers factors like tone, style, and the overall user experience. A counterfactual instance that scores high on likability is expected to be more engaging and appealing to readers.
Grammar: Evaluates the syntactical correctness and grammatical accuracy of the generated text. A counterfactual instance should follow grammatical rules, ensuring that sentences are properly constructed, punctuation is used correctly, and overall grammar is of a high standard.
The given textual metrics play a crucial role in evaluating generated text comprehensively, offering insights into various quality aspects. When combined with traditional counterfactual metrics, they provide a detailed understanding of both effectiveness and linguistic quality. Chiang and Lee (2023) demonstrated that language models (LLMs) can serve as an alternative method to assess text quality using these metrics. In this work, we use Mistral-7B-instruct Jiang et al. (2023) to evaluate the above textual quality metrics on a scale ranging from 1 to 5, as outlined in prior work Chiang and Lee (2023).
3.2 Datasets

### v2

@@ Abstract
Counterfactual text generation aims to minimally change a text, such that it is classified differently. Assessing progress in method development for counterfactual text generation is hindered by a non-uniform usage of data sets and metrics in related work. We propose CEval, a benchmark for comparing counterfactual text generation methods. CEval unifies counterfactual and text quality metrics, includes common counterfactual datasets with human annotations, standard baselines (MICE, GDBA, CREST) and the open-source language model LLAMA-2. Our experiments found no perfect method for generating counterfactual text. Methods that excel at counterfactual metrics often produce lower-quality text while LLMs with simple prompts generate high-quality text but struggle with counterfactual criteria. By making CEval available as an open-source Python library, we encourage the community to contribute additional methods and maintain consistent evaluation in future work. [^1] https://github.com/aix-group/CEval-Counterfactual-Generation-Benchmark
@@
Van Bach Nguyen University of Marburg, Germany vanbach.nguyen@uni-marburg.de                        Christin Seifert University of Marburg, Germany christin.seifert@uni-marburg.de
Jörg Schlötterer University of Marburg, Germany University of Mannheim, Germany joerg.schloetterer@uni-marburg.de
@@ 1 Introduction
The rise of deep learning and complex “black-box” models has created a critical need for interpretability. As Miller (2019) notes, explanations often involve counterfactuals to understand why event P occurred instead of Q. Ideally, these explanations show how minimal changes in an instance could lead to different outcomes. For example, to explain why the review “The film has funny moments and talented actors, but it feels long.” is negative rather than positive, a counterfactual like “The film has funny moments and talented actors, yet feels a bit long.” can be used (see Fig. 1 for more counterfactual examples generated by different methods on the same original instance). This explanation highlights specific words to change and modifications needed for a positive sentiment . It also motivates counterfactual generation, which requires modifying an instance minimally to obtain a different model prediction.
Figure 1: Examples of counterfactuals generated by different methods and human annotators that successfully flip the label from negative to positive for the same original instance.
@@
CF-GAN Robeer et al. (2021) HATESPEECH, SST-2, SNLI Fidelity, Perceptibility, Naturalness SEDC Martens and Provost (2014) PWWS+ Ren et al. (2019) Polyjuice Wu et al. (2021) TextFooler Jin et al. (2020)
CORE Dixit et al. (2022) IMDB, MNLI Diversity, Closeness, Accuracy Polyjuice Wu et al. (2021) GPT-3 Brown et al. (2020) Human-CAD
DISCO Chen et al. (2023) SNLI, WANLI Flip Score, Diversity, Accuracy Tailor Ross et al. (2022) Z-aug Wu et al. (2022) Human-CAD
Table 1: Inconsistent use of datasets, metrics, and baselines across different methods.
Besides explanations Robeer et al. (2021), the NLP community uses counterfactuals for debugging models Ross et al. (2021), data augmentation Dixit et al. (2022); Chen et al. (2023); Bhattacharjee et al. (2024), and enhancing model robustness Treviso et al. (2023); Wu et al. (2021). However, because it requires deciding where and how to change the text, with many possible modifications and a vast vocabulary. While many counterfactual generation methods for text data exist in the literature, they lack unified evaluation standards. Table 1 highlights inconsistencies in datasets, metrics, and baselines across different studies, making it difficult to compare different methods or selecting the most suitable method for specific applications. To overcome these limitations, a comprehensive benchmark to thoroughly evaluate counterfactual generation methods is necessary. A benchmark that provides standardized datasets, metrics, and baselines, enabling fair and effective comparisons, and ultimately driving progress in counterfactual generation.
This work introduces CEval, the first comprehensive benchmark for evaluating methods that modify text to change classifier predictions, including contrastive explanations, counterfactual generation, and adversarial attacks. CEval offers a robust set of metrics, incorporating established metrics from the literature alongside a novel metric we propose that captures probability changes rather than hard flip rates. This set enables the assessment of both “counterfactual-ness” (e.g., label flipping ability) and textual quality (e.g., fluency, grammar, coherence). The benchmark includes curated datasets with human annotations and a strong baseline using a large language model with a simple prompt to ensure high evaluation standards. Using CEval, we systematically review and compare state-of-the-art methods, highlighting their strengths and weaknesses in generating counterfactual text. We analyze how automatically generated counterfactuals compare to human examples, revealing gaps and opportunities for improvement. We find that counterfactual generation methods often generate text that lacks in quality compared to simple prompt-based LLMs. In contrast, while the latter typically exhibit higher text quality, they may struggle to satisfy counterfactual metrics. These insights suggest exploring combinations of both paradigms into hybrid methods as promising direction for future research. By demonstrating that an open-source LLM can serve as an alternative to a closed-source LLM in text evaluation, we make the benchmark completely open-source, thereby promoting reproducibility and facilitating further research in this domain.
@@ 2 Related Work
Terms like “counterfactual” and “contrastive” generation are often used interchangeably in literature Stepin et al. (2021) and our work adopts an inclusive definition. We define counterfactual generation as a process of generating a new instance $x^{\prime}$, from the original instance $x$, that results in a different model prediction $y^{\prime}$ with minimal changes. This definition includes counterfactual, contrastive generation, and adversarial attacks. Primarily, adversarial attacks focused on changing the label without considering text quality. Recent work like GBDA Guo et al. (2021) focuses on producing adversarial text that is more natural by adding fluency and semantic similarity losses. Hence, we include GBDA in our benchmark. Technically, counterfactual generation methods for text fall into three categories:
@@
Masking and Filling Methods (MF): These methods perform two steps: (1) identifying important words for masking by various techniques, such as selecting words with the highest gradient or training a separate rationalizer for the masking process and (2) replacing the masked words using a pretrained language model with fill-in-the-blank capability. In step (1), MICE Ross et al. (2021) and AutoCAD Wen et al. (2022) use the gradient of the classifier. DoCoGen Calderon et al. (2022) identifies all domain-specific terms by calculating a masking score for n-grams (where n3absent3$\leq$ 3≤ 3) and masks all n-grams with a masking score exceeding a threshold $\tau$. Meanwhile, CREST Treviso et al. (2023) trains SPECTRA Guerreiro and Martins (2021) as a separate rationalizer to detect which phrases or words to mask. In step (2), each of these methods fine-tunes T5 to fill in the blanks created during masking. Additionally, Polyjuice Wu et al. (2021) takes text with user-specified manual masking as input and fine-tunes a RoBERTa-based model to fill in the blanks using control codes.
@@
Conditional Distribution Methods (CD): Methods like GBDA Guo et al. (2021) and CF-GAN Robeer et al. (2021) learn a conditional distribution for counterfactuals. The counterfactuals are obtained by sampling from this distribution based on a target label.
@@
Counterfactual Generation with Large Language Models: Recently, there has been a trend towards using Large Language Models (LLMs) for counterfactual generation. Approaches like CORE Dixit et al. (2022), DISCO Chen et al. (2023) and FLARE Bhattacharjee et al. (2024) optimize prompts fed into LLMs to generate the desired counterfactuals. This trend is driven by the versatile capabilities of LLMs in various tasks Maynez et al. (2023).
@@
Despite the diverse approaches proposed in generating counterfactuals across various studies, the common objective remains to generate high-quality counterfactuals. However, previous studies employed different metrics, baselines, and datasets, as illustrated in Table 1. Therefore, given the rapid growth of approaches in this field, establishing a unified evaluation standard becomes paramount. Existing benchmarks for counterfactual generation Pawelczyk et al. (2021); Moreira et al. (2022) focus exclusively on tabular data with properties that are orthogonal to text (e.g., continuous value ranges). Hence, we introduce CEval to fill this gap and provide a standard evaluation framework specifically tailored to textual counterfactual generation. Our benchmark unifies metrics of both, counterfactual criteria and text quality assessment, including datasets with human annotations and a simple baseline from a large language model.
@@ 3 Benchmark Design
We focus on counterfactual generation for textual data, which involves editing given text with minimal modifications to produce new text that increases the probability of a predefined target label with respect to a black-box classifier. This process aims to generate a counterfactual, denoted as $x^{\prime}$, that changes the classifier’s predictions compared to the original text $x$.
Formally, given a fixed classifier $f$ and a dataset with $N$ samples $(x_{1},x_{2},\ldots,x_{N})$, $x_{i}=(z_{1},z_{2},\ldots,z_{n})$ represents a sequence of $n$ tokens. The original prediction is denoted as $f(x)=y$, while the counterfactual prediction is $y^{\prime}\neq y$. The counterfactual generation process is represented by a method $e:(z_{1},\ldots,z_{n})\mapsto(z^{\prime}_{1},\ldots,z^{\prime}_{m})$, ensuring that $f(e(x))=y^{\prime}$. The resulting counterfactual example is $x^{\prime}=(z^{\prime}_{1},\ldots,z^{\prime}_{m})$ with $m$ tokens.
@@
Predictive Probability: A counterfactual instance $x^{\prime}$ should closely produce the predefined prediction $y^{\prime}$. In other words, the counterfactual text should obtain the desired target label.
@@
Likelihood in Feature Space: A counterfactual should exhibit feature values that resemble real-world text, indicating that $x^{\prime}$ remains close to a common distribution for text. This criterion ensures that the generated text is plausible, realistic and consistent with typical language patterns.
Diversity: When an explanation is ineffective, humans can offer alternatives. Similarly, if a counterfactual is unrealistic or not actionable, it is beneficial to modify the original instance differently to provide diverse options Mothilal et al. (2020). Therefore, an effective counterfactual method should present multiple ways to change a text instance to obtain the target label. Diversity is measures for a set of counterfactual instances.
@@ 3.1.1 Counterfactual metrics
where $\mathbbm{1}$ is the indicator function.
Probability Change ($\Delta\textbf{P}$): While the flip rate offers a binary assessment of Predictive Probability, it does not capture the magnitude of change towards the desired prediction. Some instances may get really close to the target prediction but still fail to flip the label. For example, a review such as: The movie looks great but has a confusing plot and slow pacing is close to a positive label but remains negative. Consequently, its probability for the positive label should be larger than for a review like This movie is terrible, which is really negative. The Probability Change ($\Delta$P) metric captures such cases by quantifying the difference between the probability of the target label $y^{\prime}$ for the original instance $x$ and the probability of the target label for the contrasting instance $x^{\prime}$.
$\Delta P=\frac{1}{N}\sum_{i=1}^{N}\bigl{(}P(y^{\prime}_{i}\mid x^{\prime}_{i},f)-P(y^{\prime}_{i}\mid x_{i},f)\bigr{)}$
Here, $P(y\mid x,f)$ is the probability that classifier $f$ assigns to label $y$ on instance $x$.
Token Distance (TD): To measure Textual Similarity, we use the token-level Levenshtein distance $d(x,x^{\prime})$ between the original instance $x$ and the counterfactual $x^{\prime}$. This metric captures all types of text edits—insertions, deletions, and substitutions—making it ideal for evaluating minimal edits as counterfactual generation involves making these specific edits rather than completely rewriting the text. The Levenshtein distance is widely used in related work on counterfactual generation (e.g., Ross et al. (2021); Treviso et al. (2023)).
$TD=\frac{1}{N}\sum_{i=1}^{N}{d(x_{i},x^{\prime}_{i}})$
Perplexity (PPL): To evaluate whether the generated text is plausible, realistic, and follows a natural text distribution, we use perplexity from GPT-2 because of its effectiveness in capturing such distributions Radford et al. (2019). [^2] While we use GPT-2 in this study, any other LLM with strong text generation capabilities is a viable drop-in replacement.
@@
where $\log p_{\theta}(z_{i}\mid z_{<i})$ is the log-likelihood of token $z_{i}$ given the previous tokens $z_{<i}$.
Diversity (Div): We quantify diversity by measuring the token distance between pairs of generated counterfactuals. Given two counterfactuals, $x^{\prime 1}$ and $x^{\prime 2}$, for the same instance $x$, diversity is defined as the average pairwise distance between the sets of counterfactuals:
$Div=\frac{1}{N}\sum_{i=1}^{N}d(x^{\prime 1}_{i},x^{\prime 2}_{i})$
Here, $d(x^{\prime 1}_{i},x^{\prime 2}_{i})$ is the Levenshtein distance between the corresponding tokens of the two counterfactuals for the $i$-th instance.
@@ 3.1.2 Text Quality Metrics
In addition to counterfactual evaluation metrics, we measure the quality of the generated text. Text quality metrics are designed to evaluate specific aspects of texts. Following Chiang and Lee (2023); Wang et al. (2023b), key text quality metrics for comprehensive insights into text quality are: 1) Fluency – natural and readable text flow; 2) Cohesiveness – logical and coherent structure and 3) Grammar – syntactical and grammatical accuracy.
Combined with counterfactual metrics, text quality metrics provide a comprehensive view on effectiveness and linguistic quality of generated counterfactuals. Evaluating these text quality metrics usually requires human annotations, which are costly and time-consuming. Recently, Chiang and Lee (2023); Huang et al. (2023); Wang et al. (2023b) showed that LLMs, specifically GPT-3/4 and ChatGPT, can serve as an alternative to human evaluation for assessing text quality using these metrics. In this work, we use ChatGPT (gpt-3.5-turbo-0125) with a temperature of 0.2 to evaluate the above textual quality metrics on a scale from 1 to 5 following Chiang and Lee (2023); Gilardi et al. (2023).
3.2 Datasets and Classifiers